OpenStudy (anonymous):

So my teacher showed me how to solve 2 square root of 45 but i forgot how can someone help please? I have my regents tmrw morning !

zepdrix (zepdrix):

\[\Large\rm 2\sqrt{45}\]So we need to simplify this?

zepdrix (zepdrix):

You're looking for `factors` of 45 which are perfect squares.\[\Large\rm 2\cdot2=4, \qquad\qquad \text{4 is a perfect square}\]\[\Large\rm 3\cdot3=9,\qquad\qquad \text{9 is a perfect square}\]\[\Large\rm ...\]

zepdrix (zepdrix):

Let's break down 45 into factors and see if we can find a perfect square,\[\Large\rm 2\sqrt{9\cdot 5}\]Ooo we found one! 9 is a perfect square. We'll separate the 9 and 5 like this,\[\Large\rm 2\sqrt9\cdot \sqrt5\]And now we can take the square root of 9,\[\Large\rm 2\cdot3\sqrt5\]Which simplifies,\[\Large\rm 6\sqrt{5}\]

zepdrix (zepdrix):

Any confusion there? :o Hopefully I read your question correctly...

OpenStudy (anonymous):

yes it did thank you so much !